OpenStudy (yoloshroom):

Profit/Cost/Revenue Problem, bit lost. Model for profit a company makes in 1 year from producing X units of it's product is. P(x)=-x^2+600x-89750 What is the max profit the company can make? ..... Im a bit lost, test tomorrow halp Choices ate 200k 250k 275k 300k 325k

OpenStudy (yoloshroom):

@AlexandervonHumboldt2

OpenStudy (inkyvoyd):

you seek to maximize the quadratic function P(x). Recall that the maximum or minimum of a parabola occurs at the vertex of the parabola. Recall again, that for quadratic equation in form ax^2+bx+c=0, the vertex (x coordinate) occurs at x=-b/2. Apply this to your problem.

OpenStudy (yoloshroom):

mmmm -b/2 ty.

OpenStudy (yoloshroom):

so x = 300 right? so then wouldnt i plu7g 300 in for the X value?

OpenStudy (yoloshroom):

@inkyvoyd

OpenStudy (inkyvoyd):

oops I made a typo LOL it's \(\Huge \frac{-b}{2a}\)

OpenStudy (inkyvoyd):

but yes using that you'd plug in whatever you get for -b/(2a) for the x value aka P(-b/(2a)) and you would get your answer for profit.

OpenStudy (yoloshroom):

gotcha, came out the same for this one.

OpenStudy (yoloshroom):

again, need to find the max # of inches a redwood can grow per yr using g(r)=-.02r^2+r+1 same formula?

OpenStudy (inkyvoyd):

yup.; b=1, a=-.02

OpenStudy (yoloshroom):

answer comes out wrong, needs to be between 12.9 and 13.1... im confused lol

OpenStudy (yoloshroom):

@inkyvoyd

OpenStudy (yoloshroom):

wrong file lol..

OpenStudy (inkyvoyd):

?

OpenStudy (yoloshroom):

nothin, what's the final answer for this problem? i got lost.

OpenStudy (inkyvoyd):

g(r)=-.02r^2+r+1 b=1 a=-.02 -b/(2a)=1/.04=25 ... but I bet if that is not the answer than you are not giving me the full problem lol

OpenStudy (yoloshroom):

A horticulturist has determined that the number of inches a young redwood tree grows in one year is a function of the annual rainfall, r (in inches), given by g(r)=-.02r^2+r+1 What is the maximum number of inches a young redwood can grow in a year? The maximum number of inches is: (A) less than 12 (B) between 12.9 and 13.1 (C) between 13.2 and 13.6 (D) between 14.4 and 15.3 (E) between 24.7 and 25.2 answer is apparently C according to answer sheet.

OpenStudy (inkyvoyd):

unless I'm missing something I think the answer is E.

OpenStudy (yoloshroom):

answer sheet is wrong then? says C :(

OpenStudy (inkyvoyd):

might want to ask but yeah I would say so.